OpenStudy (anonymous):

what is the equation of the line in slope-intercept form that passes the point (1,3) and is perpendicular to the line y=-x+2? a.y=x-4 b. y=x+2 c.y=-x+2 d. y=-x-4

OpenStudy (anonymous):

the slope of the line is -1 plug m = -1 and x1=1, y1= 3 into the standard equation y-y1 = m(x-x1) simplify and rearrange to the required form

OpenStudy (hoblos):

the slope of the line is 1

OpenStudy (anonymous):

is it a b c or d ?

OpenStudy (hoblos):

b

OpenStudy (anonymous):

sorry slope is 1 - i misread the question

OpenStudy (anonymous):

can someone show me step by step how to do this i dont know how to do this

OpenStudy (hoblos):

since the line is perp. to y=-x+2 => slope * -1 must be = -1 => slope = 1 applying the formula: y-y1 = m(x-x1) you get y-3=1(x-1) => y=x+2

OpenStudy (anonymous):

what the formula is called ?